OpenStudy (lgbasallote):

The velocity of a particle moving along the x-axis changes from \(\text V_i\) to \(\text V_f\). For which values of \(\text V_i\) and \(\text V_f\) is the total work done on the particle negative? \[A) \; \text V_i = 2 \text{m/s}, \; \text V_f = 5\text{m/s} \quad C) \; \text V_i = -2 \text{m/s}, \text V_f = -5 \text{m/s}\] \[B) \; \text V_i = -2 \text{m/s}, \; \text V_f = 5 \text{m/s} \quad D) \; \text V_i = -5 \text{m/s}, \; \text V_f = 2 \text{m/s}\]

OpenStudy (lgbasallote):

*why* isn't it C?

OpenStudy (anonymous):

it should be c :S

OpenStudy (lgbasallote):

sadly, you and i are both wrong

OpenStudy (anonymous):

wait a moment.. \[\Delta KE = Work done\]

OpenStudy (lgbasallote):

...that means what..?

OpenStudy (lgbasallote):

\[\frac 12 mv_f ^2 - \frac 12 mv_i ^2\] right?

OpenStudy (anonymous):

\[\frac{ 1 }{ 2 } m (V_f^2-V_i^2)\]

OpenStudy (lgbasallote):

yes...that's what i said

OpenStudy (anonymous):

so if u square both.. work done is positive..

OpenStudy (anonymous):

The answer is D because in it the particle SLOWS down

OpenStudy (anonymous):

it mist be D

OpenStudy (anonymous):

2^2-5^2 = negative

OpenStudy (lgbasallote):

so it's D becaise the \(\textbf{*magnitude*}\) of the final velocity is smaller than the magnitude of the initial velocity?

OpenStudy (anonymous):

Work is quite intuitive it EnergyFinal - Energy Initial

OpenStudy (anonymous):

Yes this depends only on MAGNITUDE

OpenStudy (lgbasallote):

oh. would've been fun to have known that earlier...

OpenStudy (anonymous):

M v^2 only depends on magnitude

OpenStudy (anonymous):

Mdl ?